OpenStudy (sugarmochi):

the minute hand of a clock is 9.5cm long. How far will the tip of the hand travel in one day

OpenStudy (mrnood):

Do you knw the formula to calculate the circumference of a circle from its radius?

OpenStudy (sugarmochi):

i do know the formula for area but when i do it the answer just isnt right

OpenStudy (mrnood):

It is not area you want, it is the circumference - all the way round the circle

OpenStudy (sugarmochi):

is the circumference formula c=pi D?

OpenStudy (mrnood):

yes - or another way to put it is C= 2*pi*r (because D= 2r) So you know the radius = so calculate eth circumference. This is how far the tip travels for ONE trip around the clockface. How many times does it go around in 1 day?

OpenStudy (sugarmochi):

twice.. so you times it by 2?

OpenStudy (mrnood):

The question asks about the 'minute hand' - your answer is for the ' hour hand' How many times does the 'big hand' go around in 1 day?

OpenStudy (sugarmochi):

its 24 times you times it by 24?

OpenStudy (mrnood):

Correct So oyur answer is 2* pi * r *24

OpenStudy (sugarmochi):

so its 2 (pi) (9.5) (24) ?

OpenStudy (mrnood):

try it

OpenStudy (sugarmochi):

i did it it came out as 1432.56625

OpenStudy (sugarmochi):

is it correct?

OpenStudy (mrnood):

You need to put in units. (note the centimetres are not an 'approved' unit - but would probably be OK for this question because they gave you the radius in cm) I haven't checked your calculation - but the formula you wrote was correct

OpenStudy (mrnood):

Also - you have tooo many decimal places.....

OpenStudy (sugarmochi):

but the thing is that my sheet gives the answer and question and oyu have to do the formula your self and the answers say 68m

OpenStudy (sugarmochi):

and when i convert it to m it comes out as 14.325m

OpenStudy (mrnood):

Hmm - I have checked your calculation and it is correct. The method I gave is also correct. IF the answer you have is correct then Distance = 68m = 24* C Therefore C= 2.83m If c= 2.83m = pi*D then D = 2.83/pi = .91m= 91cm so r= 45.5cm I cannot see how that answer can be correct

OpenStudy (sugarmochi):

hmm mebe a typo

OpenStudy (sugarmochi):

but thanks soo much!!

OpenStudy (mrnood):

np